# Utility-Diversity Aware Online Batch Selection for LLM Supervised Fine-tuning

### v2

@@ B.1 Parameter Analysis
We evaluate the influence of the projected dimensions $d_{1}$, $d_{2}$, and the memory buffer size $M$ on both model performance and resource usage. As shown in Figures 6 and 7, increasing the projected dimensions and memory buffer consistently improves accuracy, which gradually saturates once $d_{1}$, $d_{2}$, and $M$ become sufficiently large. The trends of $d_{1}$ and $d_{2}$ are consistent with the Johnson-Lindenstrauss lemma (johnson1984extensions; ailon2006approximate; jin2021faster; ailon2009fast; tropp2011improved), while a sufficiently large $M$ leads to performance saturation, indicating that the stored representations are adequate to capture global sample diversity. The corresponding increases in memory usage and computation time are minor, and arise only when computing inter-sample diversity.
@@
Figure 6: Impact of low projected dimensions $d_{1}$ and $d_{2}$, and memory buffer size $M$ on model performance and additional memory usage for Qwen-2.5 on MMLU. The green curves show average accuracy (%), and the blue curves show extra peak memory consumption (GB). Performance improves with increasing projected dimensions and memory buffer size, while additional memory usage increases only slightly.
@@
B.2 Additional Ablation Studies on Buffer Update Policies, Distance Aggregation Strategies, and Random Matrix Constructions
@@
B.3 Experiments on Full SFT, Larger Batch Sizes, and Instruction-Tuned Models
@@
B.4 Additional Experiments on Tasks Requiring Long-Context Reasoning
In all previous experiments, we set the maximum sequence length to 512. To assess whether UDS can effectively capture utility and diversity for training samples that require long-context reasoning, we further conduct experiments on the MATH dataset (hendrycks2021measuring), using a maximum sequence length of 2048. We train on the MATH training split and evaluate on its test split. As shown in Table 13, UDS continues to deliver strong performance compared with other baselines, even under this more challenging long-context setting.
@@ B.4 Additional Experiments on Tasks Requiring Long-Context Reasoning
B.5 Additional Experiments on OOD Evaluation
To further assess the robustness of UDS, we conduct out-of-distribution (OOD) evaluations using GSM8K as a case study. Specifically, we train on the GSM8K training set and evaluate on its in-distribution test set as well as two OOD benchmarks: MATH500 and AMC23 (lightman2023lets). As shown in Table 14, UDS continues to exhibit strong performance across both in-distribution and OOD settings, demonstrating its robustness to distribution shifts.
@@ B.5 Additional Experiments on OOD Evaluation
B.6 Strong Correlation between the Frobenius Norm of the Logits Matrix and Sample Utility
Similar to Figures 2 and 3, we also observe a strong empirical correlation between the Frobenius norm of the logits matrix, $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$, and the change in loss, $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$, as illustrated in Figure 9.
Figure 9: Strong correlation between $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$ during the training process. Each point represents the correlation coefficient calculated from $B$ pairs of values within a single batch. We extract the result using Qwen-2.5-7B on MMLU.
B.7 Comparison with Offline Data Selection Methods
We additionally include a representative offline data selection method, FisherSFT (deb2025fishersft), which uses Fisher information gain to select a training subset, in our comparison. We conduct experiments on all four datasets using Qwen-2.5-7B as the backbone model. Since FisherSFT does not operate in the same manner as online batch selection methods, it is infeasible to fairly compare throughput, so we primarily compare the final accuracy after training. Both FisherSFT and UDS select the same fraction of data, and all other training details remain the same. The results in Table 15 show that UDS also achieves better performance than FisherSFT.
@@ B.7 Comparison with Offline Data Selection Methods
Appendix C Theoretical Analysis
C.1 Relationship Between Frobenius Norm and Nuclear Norm
@@
C.2 Optimization Analysis using Adam
When using the Adam (kingma2014adam) as optimizer, the parameter update at step $t$ is
@@ C.2 Optimization Analysis using Adam
C.3 Distance Preservation under Low-dimensional Projection
@@
Let $\bm{u}_{1},\dots,\bm{u}_{n}\in\mathbb{R}^{NV}$ denote $n$ vectors of interest. Consider all pairwise differences $\bm{u}_{i}-\bm{u}_{j}$, $1\leq i<j\leq n$; there are at most n(n1)/2n(n-1)/2 such vectors. By standard results on Fast JL transforms (SRFT/FJLT) (ailon2006approximate; jin2021faster; ailon2009fast; tropp2011improved), for any $0<\epsilon<1$ and a fixed set of $n$ vectors in $\mathbb{R}^{NV}$, a random SRFT $\bm{\Gamma}\in\mathbb{R}^{d_{1}d_{2}\times NV}$ preserves the $\ell_{2}$ norms up to $(1\pm\epsilon)$ with high probability provided
@@
Appendix D Detailed Experimental Setting
D.1 Datasets
We conduct our experiments on the following datasets, with important information listed in Table 16. Brief introduction for each dataset are provided in the following:
  • MMLU (hendryckstest2021) is a benchmark designed to assess general knowledge understanding, consisting of multiple-choice questions from various branches of knowledge. It covers 57 tasks including elementary mathematics, US history, computer science, law, and more.
  • ScienceQA (lu2022learn) is a multimodal science question answering dataset collected from elementary and high school science curricula, including multiple-choice problems that align with California Common Core Content Standards. The questions in dataset are sourced from open resources managed by IXL Learning, an online learning platform curated by experts in the field of K-12 education. We extract textual parts within it following li2024mixlora.
  • GSM8K (cobbe2021gsm8k) is a dataset of 8.5K high quality linguistically diverse grade school math word problems, each requiring multi-step arithmetic reasoning. Answers are provided in a step-by-step explanation format.
  • CodeAlpaca-20k (codealpaca) is an instruction-following dataset consisting of 20k examples generated to align code-related prompts with helpful outputs. It is synthetically constructed to enhance code instruction tuning.
  • HumanEval (chen2021evaluating) is a code generation benchmark including 164 Python programming problems. The dataset was handwritten to ensure not to be included in the training set of code generation models.
Table 16: Details of MMLU, ScienceQA, GSM8K, CodeAlpaca and HumanEval Datasets. We list the number of training and testing samples and task types for the following datasets used in our experiments.
@@ D.1 Datasets
MMLU (hendryckstest2021) $99842$ $14042$ Multiple Choice
ScienceQA (lu2022learn) $12726$ $4241$ Multiple Choice
GSM8K (cobbe2021gsm8k) $7473$ $1319$ Math Problems
CodeAlpaca-20k (codealpaca) $20022$ - Code Instruction
HumanEval (chen2021evaluating) - $164$ Code Generation
D.2 Training Configuration
@@
Appendix E Limitations and Future Work
@@
One potential limitation of our framework lies in the computation of the nuclear norm. While it can be obtained directly from the singular values of the logits matrix without incurring additional expensive backpropagation, the SVD of large matrices may still introduce non-negligible overhead in practice, particularly for long sequences with large vocabulary sizes. We also explored approximate methods such as randomized SVD to accelerate the computation, but observed that the loss in precision often degrades the overall performance. Another limitation is the difficulty of providing a rigorous theoretical proof for the linear correlation between the nuclear norm of the logits matrix and the resulting loss reduction or effective matrix rank.
@@
An important future direction is to explore more efficient and accurate estimators of the nuclear norm. Potential avenues include exploiting structured low-rank approximations, leveraging iterative or block-wise SVD techniques, or designing surrogate scoring functions that preserve the optimization and diversity signals of the nuclear norm while being cheaper to compute. Such improvements would further enhance the scalability of our framework to larger models and longer input sequences. Furthermore, techniques in uncertainty estimation or active learning may be other promising directions for online batch selection (ren2021survey; loquercio2020general; peng2024energy).
Appendix F LLM Usage Declaration
In preparing this manuscript, we used a large language model solely as a language assistance tool. Specifically, the LLM was employed to polish the phrasing of certain paragraphs and to improve clarity and readability of the text. All technical ideas, derivations, experiments, and analyses were conceived, implemented, and validated entirely by the authors. The LLM was not used for generating research ideas, designing experiments, or producing novel scientific content. The authors take full responsibility for all contents of the paper.

### v3

@@ C.1 Parameter Analysis
We evaluate the influence of the projected dimensions $d_{1}$, $d_{2}$, and the memory buffer size $M$ on both model performance and resource usage. As shown in Figures 4 and 6, increasing the projected dimensions and memory buffer consistently improves accuracy, which gradually saturates once $d_{1}$, $d_{2}$, and $M$ become sufficiently large. The trends of $d_{1}$ and $d_{2}$ are consistent with the Johnson-Lindenstrauss lemma (Johnson et al., 1984; Ailon and Chazelle, 2006; Jin et al., 2021; Ailon and Chazelle, 2009; Tropp, 2011), while a sufficiently large $M$ leads to performance saturation, indicating that the stored representations are adequate to capture global sample diversity. The corresponding increases in memory usage and computation time are minor, and arise only when computing inter-sample diversity.
@@
C.2 Additional Ablation Studies on Buffer Update Policies, Distance Aggregation Strategies, and Random Matrix Constructions
@@
C.3 Experiments on Full SFT, Larger Batch Sizes, and Instruction-Tuned Models
@@
C.4 Additional Experiments on Tasks Requiring Long-Context Reasoning
In all previous experiments, we set the maximum sequence length to 512. To assess whether UDS can effectively capture utility and diversity for training samples that require long-context reasoning, we further conduct experiments on the MATH dataset (Hendrycks et al., 2021b), using a maximum sequence length of 2048. We train on the MATH training split and evaluate on its test split. As shown in Table 13, UDS continues to deliver strong performance compared with other baselines, even under this more challenging long-context setting.
@@ C.4 Additional Experiments on Tasks Requiring Long-Context Reasoning
C.5 Additional Experiments on OOD Evaluation
To further assess the robustness of UDS, we conduct out-of-distribution (OOD) evaluations using GSM8K as a case study. Specifically, we train on the GSM8K training set and evaluate on its in-distribution test set as well as two OOD benchmarks: MATH500 and AMC23 (Lightman et al., 2023). As shown in Table 14, UDS continues to exhibit strong performance across both in-distribution and OOD settings, demonstrating its robustness to distribution shifts.
@@ C.5 Additional Experiments on OOD Evaluation
C.6 Comparison with Offline Data Selection Methods
We additionally include a representative offline data selection method, FisherSFT (Deb et al., 2025), which uses Fisher information gain to select a training subset, in our comparison. We conduct experiments on all four datasets using Qwen-2.5-7B as the backbone model. Since FisherSFT does not operate in the same manner as online batch selection methods, it is infeasible to fairly compare throughput, so we primarily compare the final accuracy after training. Both FisherSFT and UDS select the same fraction of data, and all other training details remain the same. The results in Table 15 show that UDS also achieves better performance than FisherSFT.
@@ C.6 Comparison with Offline Data Selection Methods
C.7 Analysis on Considering Within-Batch Diversity
We conducted an additional experiment showing that incorporating diversity calculation within the batch yields no notable difference in accuracy. This is because the memory buffer size is significantly larger than the batch size; thus, it is less likely that a similar sample appears in the buffer while being absent from the candidate batch. This comparison follows the experimental setting in Table 3 using Qwen-2.5-7B.
Table 16: Accuracy Comparison With and Without Within-batch Diversity. We report the average accuracy using Qwen-2.5-7B.
Method MMLU ScienceQA GSM8K HumanEval
[…]
Appendix D Theoretical Analysis
D.1 Relationship Between Frobenius Norm and Nuclear Norm
@@
D.2 Optimization Analysis using Adam
When using the Adam (Kingma and Ba, 2014) as optimizer, the parameter update at step $t$ is
@@ D.2 Optimization Analysis using Adam
D.3 Distance Preservation under Low-dimensional Projection
@@
Let $\bm{u}_{1},\dots,\bm{u}_{n}\in\mathbb{R}^{NV}$ denote $n$ vectors of interest. Consider all pairwise differences $\bm{u}_{i}-\bm{u}_{j}$, $1\leq i<j\leq n$; there are at most n(n1)/2n(n-1)/2 such vectors. By standard results on Fast JL transforms (SRFT/FJLT) (Ailon and Chazelle, 2006; Jin et al., 2021; Ailon and Chazelle, 2009; Tropp, 2011), for any $0<\epsilon<1$ and a fixed set of $n$ vectors in $\mathbb{R}^{NV}$, a random SRFT $\bm{\Gamma}\in\mathbb{R}^{d_{1}d_{2}\times NV}$ preserves the $\ell_{2}$ norms up to $(1\pm\epsilon)$ with high probability provided
@@
Appendix E Detailed Experimental Setting
E.1 Datasets
We conduct our experiments on the following datasets, with important information listed in Table 17. Typically, for each reported result, we train the model on a single training set and evaluate it on a single test set. Specifically, for MMLU, models were trained on MMLU’s auxiliary dataset and tested on its test set; for GSM8K, models were trained on GSM8K’s training set and tested on its test set; for ScienceQA, models were trained on its training set and tested on its test set; for HumanEval, models were trained on CodeAlpaca and evaluated on the HumanEval benchmark; and for MATH, models were trained on MATH’s training set and tested on its test set. Specifically, for the OOD experiments, we evaluate the model trained on GSM8K directly on the MATH500 and AMC23 datasets. Brief introduction for each dataset are provided in the following:
  • MMLU (Hendrycks et al., 2021a) is a benchmark designed to assess general knowledge understanding, consisting of multiple-choice questions from various branches of knowledge. It covers 57 tasks including elementary mathematics, US history, computer science, law, and more.
  • ScienceQA (Lu et al., 2022) is a multimodal science question answering dataset collected from elementary and high school science curricula, including multiple-choice problems that align with California Common Core Content Standards. The questions in dataset are sourced from open resources managed by IXL Learning, an online learning platform curated by experts in the field of K-12 education. We extract textual parts within it following (Li et al., 2024).
  • GSM8K (Cobbe et al., 2021) is a dataset of 8.5K high quality linguistically diverse grade school math word problems, each requiring multi-step arithmetic reasoning. Answers are provided in a step-by-step explanation format.
  • CodeAlpaca-20k (Chaudhary, 2023) is an instruction-following dataset consisting of 20k examples generated to align code-related prompts with helpful outputs. It is synthetically constructed to enhance code instruction tuning.
  • HumanEval (Chen et al., 2021) is a code generation benchmark including 164 Python programming problems. The dataset was handwritten to ensure not to be included in the training set of code generation models.
  • MATH (Hendrycks et al., 2021b) consists of 12.5K challenging competition mathematics problems from high school math competitions. It covers diverse subjects including algebra, geometry, counting, and calculus, requiring complex heuristic reasoning to solve.
  • MATH500 (Lightman et al., 2023) is a curated subset of 500 examples from the MATH dataset’s test split. It is widely used to evaluate the mathematical reasoning capabilities of LLMs in a more computationally efficient manner while maintaining the difficulty distribution of the original benchmark.
  • AMC23 (Lightman et al., 2023) comprises problems from the 2023 American Mathematics Competitions (AMC 12). It serves as an out-of-distribution (OOD) benchmark to assess the model’s ability to generalize to recent mathematical problems that were not seen during the training phase.
Table 17: Details of MMLU, ScienceQA, GSM8K, CodeAlpaca, HumanEval, MATH, MATH500 and AMC23 Datasets. We list the number of training and testing samples and task types for the following datasets used in our experiments.
@@ E.1 Datasets
MMLU (Hendrycks et al., 2021a) $99842$ $14042$ Multiple Choice
ScienceQA (Lu et al., 2022) $12726$ $4241$ Multiple Choice
GSM8K (Cobbe et al., 2021) $7473$ $1319$ Math Problems
CodeAlpaca-20k (Chaudhary, 2023) $20022$ - Code Instruction
HumanEval (Chen et al., 2021) - $164$ Code Generation
MATH (Hendrycks et al., 2021b) $7500$ $5000$ Math Problems
MATH500 (Lightman et al., 2023) - $500$ Math Problems
AMC23 (Lightman et al., 2023) - $40$ Math Problems
E.2 Training Configuration
@@
Appendix F Limitations and Future Work
@@
One potential limitation of our framework lies in the computation of the nuclear norm. While it can be obtained directly from the singular values of the logits matrix without incurring additional expensive backpropagation, the SVD of large matrices may still introduce non-negligible overhead in practice, particularly for long sequences with large vocabulary sizes. We also explored approximate methods such as randomized SVD to accelerate the computation, but observed that the loss in precision often degrades the overall performance. Another limitation is the difficulty of providing a rigorous theoretical proof for the linear correlation between the nuclear norm of the logits matrix and the resulting loss reduction or effective matrix rank given the complexity of the large model.
@@
An important future direction is to explore more efficient and accurate estimators of the nuclear norm. Potential avenues include exploiting structured low-rank approximations, leveraging iterative or block-wise SVD techniques, or designing surrogate scoring functions that preserve the optimization and diversity signals of the nuclear norm while being cheaper to compute. Such improvements would further enhance the scalability of our framework to larger models and longer input sequences.